AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 316     Entry time: Sun Feb 27 23:17:33 2022
Author: JM, OF 
Subject: Current status 

00:20

Temperature setpoint 115 K

Rutherford rate appears fairly stable. Fluctuations from a max peak of 10^2 to a minimum of ~40 peak hieght. Fluctuates remain lower than at 95 K.

Temperatures OK

System checks OK - white rabbit attached.

Leakage current ~ 3.236 uA

Data rate ~1177 kb/sec

Merger OK

Ring vacuum levels appear to have stablised from previous large fluctuations. Avg ring pressure decreased to 3.8E-10 mbar from 7.5E-10 mbar at start of the shift. (attachment 9)

00:51

Rutherford rate still stable. Fluctuations have reduced slightly. Min peak height now ~60 counts, max peak ~10^2 counts. (attachment 10)

Ring vacuum avg = 3.2 E-10 mbar.

E1 pump current ~ 2.4 A

Attachment 1: 2022-02-28_00-19-06-Rate.png  155 kB  | Hide | Hide all
2022-02-28_00-19-06-Rate.png
Attachment 2: 2022-02-28_00-22-09-wr.png  127 kB  | Hide | Hide all
2022-02-28_00-22-09-wr.png
Attachment 3: 2022-02-28_00-22-37-good-events.png  103 kB  | Hide | Hide all
2022-02-28_00-22-37-good-events.png
Attachment 4: 2022-02-28_00-23-09-ADCitems.png  103 kB  | Hide | Hide all
2022-02-28_00-23-09-ADCitems.png
Attachment 5: 2022-02-28_00-23-35-correlation.png  102 kB  | Hide | Hide all
2022-02-28_00-23-35-correlation.png
Attachment 6: 2022-02-28_00-25-39-ExTotal.png  169 kB  | Hide | Hide all
2022-02-28_00-25-39-ExTotal.png
Attachment 7: 2022-02-28_00-26-42-xyTotal.png  268 kB  | Hide | Hide all
2022-02-28_00-26-42-xyTotal.png
Attachment 8: 20220228_001956.jpg  863 kB  | Hide | Hide all
20220228_001956.jpg
Attachment 9: 20220228_003157.jpg  1.586 MB  | Hide | Hide all
20220228_003157.jpg
Attachment 10: 2022-02-28_00-51-30-rates.png  156 kB  | Hide | Hide all
2022-02-28_00-51-30-rates.png
ELOG V3.1.4-unknown